train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and train B is traveling at 80 miles per hour. Train A passes a station at 6:10 P.M. If train B passes the same station at 6:25 P.M., at what time will train B catch up to train A?
Answer by Alan3354(69443)

train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and train B is traveling at 80 miles per hour. Train A passes a station at 6:10 P.M. If train B passes the same station at 6:25 P.M., at what time will train B catch up to train A?
--------------
In 1/4 hour (625 - 610) train A goes 15 miles (60 * 1/4)
Train B gains on A at 20 mph (80 - 60)
15 miles/20 mph = 3/4 hour = 45 minutes
6:25 + 0:45 = 7:10 PM